OpenStudy (anonymous):

Find the 5th term of (x-1)^7

OpenStudy (alekos):

use the binomial theorem

OpenStudy (anonymous):

yes, I posted the steps to solving with the binomial theorem in the last question you asked. I can help guide you through it if you'd like

OpenStudy (anonymous):

I just need help on the exponents for (x) and (-1) So far I have 7 C 4 (x) (-1)

OpenStudy (nincompoop):

\[(a+b)^n= \sum_{i=0}^{n}\left(\begin{matrix}n \\ i\end{matrix}\right)n^{n-i}b^i\]

OpenStudy (nincompoop):

review Pascal's Triangle so you will see the pattern

OpenStudy (anonymous):

Okie thanks